OpenStudy (anonymous):

50 is 125% of what number? 12.5 40 62.5 112.5

OpenStudy (anonymous):

50 is 125% of what number. Here "is" means "=" "of" means "multiple" "what number" means "x" (soem number) So first translate it and then it would be easy :) 50 = 125% of x 50 = 125/100 * x 50 = 5/4 * x x = 50 * 4/5 x = 50/5 * 4 = 10 * 4 = 40 x = 40 So, 50 is 125% of 40 :)